OpenStudy (danmac0710):

I have a graph of a function with equation; y=ln(ax+b). It has a vertical asymptote at x=1 and x-intercept at (4,0). Can you remind me of an efficient way to find parameters 'a' and 'b' please? I was thinking of relating it to transformations but it seems too long a method. Thanks.

OpenStudy (dumbcow):

ok ln[a(x-h)] shifts graph "h" units left/right where x = h is vertical asymptote we are given that h=1 --> y = ln[a(x-1)] now plug in x-intercept (4,0) 0 = ln(3a) ---> 3a = 1 ---> a = 1/3 y = ln (1/3x - 1/3) a = 1/3 b = -1/3

OpenStudy (danmac0710):

great stuff - thanks mate! You explained it really nicely.

OpenStudy (anonymous):

Vertical retricemtote means undefined. You know that ln(0) = undefined right? and ln(ax+b). In other words a+b = 0 Now the other equatioin, you know that ln(1) = 0 so it would generate (4,0) so ln(1) = ln(ax+b), conclude that ax+b = 1 . Now plug in x=4 4a +b = 1 Now you have 2 equations: 4a + b = 1 a+ b= 0 b= -a substitute 4a -a =1 a = 1/3 a+ b = 0 1/3 + b = 0 b = -1/3

OpenStudy (danmac0710):

that was a good response too. Thank you too Study100. I know this stuff but I am out of practice so your help is much appreciated.